OpenStudy (anonymous):

1. One way to show that a statement is not a good definition is to find a ____. (1 point) converse conditional biconditional counterexample

OpenStudy (aroub):

I think its counterexample :)

jhonyy9 (jhonyy9):

yes counterexample is right sure

OpenStudy (aroub):

^_^

OpenStudy (anonymous):

1. One way to show that a statement is not a good definition is to find a ____. (1 point) Unselected answer (0 pts) converse Unselected answer (0 pts) conditional Unselected answer (0 pts) biconditional Correct answer (1 pt) counterexample 1 /1 point 2. Which choice shows a true conditional with a correctly identified hypothesis and conclusion? (1 point) Incorrect Answer (0 pts) If next month is January, then this month is the last of the year. Hypothesis: This month is the last of the year. Conclusion: Next month is January. Unselected answer (0 pts) Next month is February, so this month is the last month of the year. Hypothesis: Next month is February. Conclusion: This month is the last of the year. Unselected answer (0 pts) Next month is February, so this month is the last month of the year. Hypothesis: This month is the last of the year. Conclusion: Next month is February. Unselected answer (1 pt) If next month is January, then this month is the last of the year. Hypothesis: Next month is January. Conclusion: This month is the last of the year. 0 /1 point 3. What conjecture can you make about the sum of the first 10 positive even numbers? Geometry.02.Quiz.Q5 (1 point) Unselected answer (0 pts) The sum is 9 · 10. Unselected answer (0 pts) The sum is 10 · 10. Unselected answer (0 pts) The sum is 11 · 12. Correct answer (1 pt) The sum is 10 · 11. 1 /1 point 4. Is the following definition of perpendicular reversible? If yes, write it as a true biconditional. Two lines that intersect at right angles are perpendicular. (1 point) Unselected answer (0 pts) The statement is not reversible. Unselected answer (0 pts) Yes; if two lines intersect at right angles, then they are perpendicular. Unselected answer (0 pts) Yes; if two lines are perpendicular, then they intersect at right angles. Correct answer (1 pt) Yes; two lines intersect at right angles if (and only if) they are perpendicular. 1 /1 point 5. Which biconditional is not a good definition? (1 point) Incorrect Answer (0 pts) A whole number is odd if and only if the number is not divisible by 2. Unselected answer (0 pts) An angle is straight if and only if its measure is 180°. Unselected answer (0 pts) A whole number is even if and only if it is divisible by 2. Unselected answer (1 pt) A ray is a bisector of an angle if and only if it splits the angle into two angles. 0 /1 point 6. Use the Law of Detachment to draw a conclusion from the two given statements. If not possible, write not possible. If I miss a basketball practice, then I cannot play in the upcoming game. I cannot play in the upcoming game. (1 point) Unselected answer (0 pts) I missed basketball practice. Unselected answer (0 pts) I did not miss basketball practice. Unselected answer (0 pts) If I miss the upcoming basketball game, then I missed practice. Correct answer (1 pt) not possible 1 /1 point 7. Name the Property of Congruence that justifies this statement: Geometry.02.Quiz.Q12 (1 point) Unselected answer (1 pt) Transitive Property Unselected answer (0 pts) Symmetric Property Incorrect Answer (0 pts) Reflexive Property Unselected answer (0 pts) none of these 0 /1 point 8. What is a counterexample for the conjecture? Conjecture: Any number that is divisible by 4 is also divisible by 8. (1 point) Unselected answer (0 pts) 24 Unselected answer (0 pts) 40 Correct answer (1 pt) 12 Unselected answer (0 pts) 26 1 /1 point 9. Use the Law of Syllogism to draw a conclusion from the two given statements. If two lines intersect and form right angles, then the lines are perpendicular. If two lines are perpendicular, then they intersect and form 90° angles. (1 point) Unselected answer (0 pts) The lines intersect and form 90° angles. Unselected answer (0 pts) If two lines do not intersect and form 90° angles, then they do not form right angles. Unselected answer (0 pts) The lines are perpendicular. Correct answer (1 pt) If two lines intersect and form right angles, then they intersect and form 90° angles. 1 /1 point 10. For the following true conditional statement, write the converse. If the converse is also true, combine the statements as a biconditional. If x = 3, then x2 = 9. (1 point) Unselected answer (0 pts) If x2 = 9, then x = 3. True; x2 = 9 if and only if x = 3. Unselected answer (0 pts) If x2 = 3, then x = 9. False. Incorrect Answer (0 pts) If x2 = 9, then x = 3. True; x = 3 if and only if x2 = 9. Unselected answer (1 pt) If x2 = 9, then x = 3. False. 0 /1 point 11. Laisha’s Internet Services designs websites and recently began a weekly advertising campaign. Laisha noticed an increase in her customers over a period of five consecutive weeks. Based on the pattern shown in the graph, make a conjecture about the number of customers Laisha will have in the seventh week. Geometry.02.Quiz.Q18 (1 point) Unselected answer (0 pts) Laisha will have 7 customers. Unselected answer (0 pts) Laisha will have 9 customers. Unselected answer (0 pts) Laisha will have 11 customers. Correct answer (1 pt) Laisha will have 13 customers. 1 /1 point 12. Use the given property to complete the statement. Multiplication Property of Equality If 6x ÷ 8 = 40, then _____ . (1 point) Unselected answer (0 pts) 6x ÷ 8 = 320 Unselected answer (0 pts) 40 = 6x · 8 Correct answer (1 pt) 6x = 320 Unselected answer (0 pts) 40 = 6x ÷ 8 1 /1 point 13. Which statement is an example of the Addition Property of Equality? (1 point) Unselected answer (0 pts) If p = q then p · s = q · s. Correct answer (1 pt) If p = q then p + s = q + s. Unselected answer (0 pts) If p = q then p – s = q – s. Unselected answer (0 pts) p = q 1 /1 point 14. Find the values of x and y. (1 point) Unselected answer (0 pts) x = 129, y = 51 Unselected answer (0 pts) x = 51, y = 129 Correct answer (1 pt) x = 66, y = 17 Unselected answer (0 pts) x = 17, y = 66 1 /1 point 15. What is the value of x? (1 point) Correct answer (1 pt) 61 Unselected answer (0 pts) 37 Unselected answer (0 pts) 85 Unselected answer (0 pts) 122 1 /1 point 16. Name the Property of Equality that justifies this statement: If p = q, then q = p. (1 point) Unselected answer (0 pts) Reflexive Property Unselected answer (0 pts) Multiplication Property Unselected answer (1 pt) Symmetric Property Incorrect Answer (0 pts) Subtraction Property 0 /1 point Fill in the Blank 17. GEO unit 2 test Q9 image (1 point) Correct answer (1 pt) A (1 pt) A.